### v3

@@ Abstract
[ABSTRACT]Recently, Multimodal Large Language Model (MLLM) represented by GPT-4V has been a new rising research hotspot, which uses powerful Large Language Models (LLMs) as a brain to perform multimodal tasks. The surprising emergent capabilities of MLLM, such as writing stories based on images and Optical Character Recognition (OCR)-free math reasoning, are rare in traditional multimodal methods, suggesting a potential path to artificial general intelligence. To this end, both academia and industry have endeavored to develop MLLMs that can compete with or even outperform GPT-4V, pushing the limit of research at a surprising speed. In this paper, we aim to trace and summarize the recent progress of MLLMs. First of all, we present the basic formulation of MLLM and delineate its related concepts, including architecture, training strategy and data, as well as evaluation. Then, we introduce research topics about how MLLMs can be extended to support more granularity, modalities, languages, and scenarios. We continue with multimodal hallucination and extended techniques, including Multimodal ICL (M-ICL), Multimodal CoT (M-CoT), and LLM-Aided Visual Reasoning (LAVR). To conclude the paper, we discuss existing challenges and point out promising research directions.
doi:
10.1093/nsr/nwae403
keywords:
multimodal large language model, vision language model, large language model
@@ 1 INTRODUCTION
Recent years have seen the remarkable progress of LLMs [1, 2]. By scaling up data size and model size, these LLMs raise extraordinary emergent abilities, typically including instruction following  [3], In-Context Learning (ICL)  [4], and Chain of Thought (CoT)  [5]. Although LLMs have demonstrated surprising zero/few-shot reasoning performance on most Natural Language Processing (NLP) tasks [6] and even complex real-life applications [7, 8, 9], they are inherently “blind” to vision since they can only understand discrete text. At the same time, Large Vision Models (LVMs) can see clearly [10, 11], but commonly lag in reasoning.
@@
In light of this complementarity, LLM and LVM run towards each other, leading to the new field of Multimodal Large Language Model (MLLM). Formally, it refers to the LLM-based model with the ability to receive, reason, and output with multimodal information. Prior to MLLM, there have been a lot of works devoted to multimodality, which can be divided into discriminative [12, 13] and generative [14, 15] paradigms. CLIP [12], as a representative of the former, projects visual and textual information into a unified representation space, building a bridge for downstream multimodal tasks. In contrast, OFA [14] is a representative of the latter, which unifies multimodal tasks in a sequence-to-sequence manner. MLLM can be classified as the latter according to the sequence operation, but it manifests two distinct traits compared with its traditional counterparts: (1) MLLM is based on LLM with billion-scale parameters, which is not available in previous models. (2) MLLM uses new training paradigms to unleash its full potential, such as using multimodal instruction tuning [16] to encourage the model to follow new instructions. Armed with the two traits, MLLM exhibits new capabilities, such as writing website code based on images [17], understanding the deep meaning of a meme [18], and OCR-free math reasoning [19].
@@
Figure 1: A timeline of representative MLLMs. We are witnessing rapid growth in this field. More works can be found on our released GitHub page, which is updated daily.
@@
Ever since the release of GPT-4 [20], there has been a research frenzy over MLLMs because of the amazing multimodal examples it shows. Rapid development is fueled by efforts from both academia and industry. Preliminary research on MLLMs focuses on text content generation grounded in text prompts and image [16]/video [21, 22]/audio [23]. Subsequent works have expanded the capabilities or the usage scenarios, including: (1) Better granularity support. Finer control on user prompts is developed to support specifying regions through boxes [24] or a certain object through a click [25]. (2) Enhanced support on input and output modalities [26, 27], such as image, video, audio, and point cloud. (3) Improved language support. Efforts have been made to extend the success of MLLMs to other languages (e.g. Chinese) with relatively limited training corpus [28]. (4) Extension to more realms and usage scenarios. Some studies transfer the strong capabilities of MLLMs to other domains, such as medical image understanding [29] and document parsing [30]. Moreover, multimodal agents are developed to assist in real-world interaction, e.g. embodied agents [31] and GUI agents [32]. An MLLM timeline is illustrated in Fig. 1.
@@
The following parts of the survey are structured as such: the survey starts with a comprehensive review of the essential aspects of MLLMs, including (1) Mainstream architecture; (2) A full recipe of training strategy and data; (3) Common practices of performance evaluation. Then, we delve into a deeper discussion on some important topics about MLLMs, each focusing on a main problem: (1) What aspects can be further improved or extended? (2) How to relieve the multimodal hallucination issue? The survey continues with the introduction of three key techniques, each specialized in a specific scenario: M-ICL is an effective technique commonly used at the inference stage to boost few-shot performance. Another important technique is M-CoT, which is typically used in complex reasoning tasks. Afterward, we delineate general ideas for developing LLM-based systems to solve composite reasoning tasks or to address common user queries. We finish our survey with a summary and potential research directions.
@@ 2 ARCHITECTURE
Table 1: A summary of commonly used image encoders.
@@ 2.1 Modality encoder
The encoders compress raw information, such as images or audio, into a more compact representation. Rather than training from scratch, a common approach is to use a pre-trained encoder that has been aligned to other modalities. For example, CLIP [12] incorporates a visual encoder semantically aligned with the text through large-scale pre-training on image-text pairs. Therefore, it is more practical to utilize such pre-aligned encoders to align with LLMs through alignment pre-training.
Commonly used image encoders are summarized in Table 1. Apart from vanilla CLIP image encoders [12], some works also explore using other variants. For example, MiniGPT-4 [17] adopts an EVA-CLIP [36] (ViT-G/14) encoder, which is trained with improved training techniques. Osprey [25] introduces a convolution-based ConvNext-L encoder [33] to utilize higher resolution and multi-level features. Some works also explore encoder-free architecture. For instance, the image patches of Fuyu-8b [37] are directly projected before sending to LLMs. With this design, the model naturally supports flexible input image resolution.
When choosing encoders, one often considers factors like resolution, parameter size, and pretraining corpus. Notably, many works have empirically verified that using higher resolution can achieve remarkable performance gains [28, 38]. The approaches for scaling up input resolution can be categorized into direct scaling and patch-division methods. The direct scaling way inputs images of higher resolutions to the encoder, which often involves further tuning the encoder [28] or replacing a pre-trained encoder with higher resolution [39]. Similarly, CogAgent [32] uses a dual-encoder mechanism, where two encoders process high and low-resolution images, respectively. High-resolution features are injected into the low-resolution branch through cross-attention. Patch-division methods cut a high-resolution image into patches and reuse the low-resolution encoder. For example, Monkey [38] and SPHINX [40] divide a large image into smaller patches and send sub-images together with a downsampled high-resolution image to the image encoder, where the sub-images and the low-resolution image capture local and global features, respectively. In contrast, parameter size and training data composition are of less importance compared with input resolution, found by empirical studies [41].
Similar encoders are also available for other modalities. For example, Pengi [23] uses CLAP [42] model as the audio encoder. ImageBind-LLM [26] uses the ImageBind [43] encoder, which supports encoding image, text, audio, depth, thermal, and Inertial Measurement Unit (IMU) data. Equipped with the strong encoder, ImageBind-LLM can respond to the input of various modalities.
Table 2: A summary of commonly used open-sourced LLMs. en, zh, fr, and de stand for English, Chinese, French, and German, respectively.
@@ 2.2 Pre-trained LLM
We summarize the commonly used and publicly available LLMs in Table 2. Notably, most LLMs fall in the causal decoder category, following GPT-3 [4]. Among them, Flan-T5 [44] series are relatively early LLMs used in works like BLIP-2 [50] and InstructBLIP [51]. LLaMA series [45] and Vicuna family [46] are representative open-sourced LLMs that have attracted much academic attention. Since the two LLMs are mainly pre-trained on English corpus, they are limited in multi-language support, such as Chinese. In contrast, Qwen [48] is a bilingual LLM with Chinese and English support.
It should be noted that scaling up the parameter size of LLMs also brings additional gains, similar to the case of increasing input resolution. Specifically, Liu et al. [39, 52] find that simply scaling up LLM from 7B to 13B brings comprehensive improvement on various benchmarks. Furthermore, when using a 34B LLM, the model shows emergent zero-shot Chinese capability, given that only English multimodal data are used during training. Lu et al. [53] see a similar phenomenon by scaling up LLMs from 13B to 35B and 65B/70B, where the larger model size brings consistent gains on benchmarks specifically designed for MLLMs. Some works instead use smaller LLMs to facilitate deployment on mobile devices. For example, MobileVLM series [54] use downscaled LLaMA [45] to enable efficient inference on mobile processors.
Recently, explorations of Mixture of Experts (MoE) architecture for LLMs have garnered rising attention [55]. Compared with dense models, the sparse architecture enables scaling up total parameter size without increasing computational cost, by selective activation of the parameters. Empirically, MM1 [41] and MoE-LLaVA [56] find that MoE implementation achieves better performance than the dense counterpart on almost all the benchmarks.
@@ 2.3 Modality interface
Since LLMs can only perceive text, bridging the gap between natural language and other modalities is necessary. Nevertheless, it would be costly to train from scratch a large multimodal model in an end-to-end manner. A more practical way is to introduce a learnable connector between the pre-trained visual encoder and LLM. The other approach is to translate images into languages with the help of expert models, and then send the language to LLM.
Learnable Connector. It is responsible for bridging the gap between different modalities. Specifically, the module projects information into the space that LLM can understand efficiently. Based on how multimodal information is fused, there are broadly two ways to implement such interfaces, i.e. token-level and feature-level fusion for different modalities.
For token-level fusion, features output from encoders are transformed into tokens and concatenated with text tokens before being sent into LLMs. A common solution is to leverage a group of learnable query tokens to extract information in a query-based manner [57], which first has been implemented in BLIP-2 [50], and subsequently inherited by a variety of work [51, 22]. Such Q-Former-style approaches compress visual tokens into a smaller number of representation vectors. In contrast, some methods simply use a MLP-based interface to bridge the modality gap [16]. For example, LLaVA series adopts an MLP [16, 39] to project visual tokens and align the feature dimension with word embeddings. BLIVA [58] adopts an ensemble of MLP-based and Q-Former-based connector to enhance performance in text-rich scenarios.
As another line, feature-level fusion inserts extra modules that enable deep interaction and fusion between text features and visual features. For example, Flamingo [59] inserts extra cross-attention layers between frozen Transformer layers of LLMs, thereby augmenting language features with external visual cues. Similarly, CogVLM [60] plugs in a visual expert module in each Transformer layer to enable dual interaction and fusion between vision and language features. For better performance, the QKV weight matrix of the introduced module is initialized from the pre-trained LLM. Likewise, LLaMA-Adapter [61] introduces learnable prompts into Transformer layers. These prompts are first embedded with visual knowledge and then concatenated with text features as prefixes.
@@
Expert Model. Apart from the learnable interface, using expert models, such as an image captioning model, is also a feasible way to bridge the modality gap [63]. The basic idea is to convert multimodal inputs into languages without training. In this way, LLMs can understand multimodality by the converted languages. For example, VideoChat-Text [21] uses pre-trained vision models to extract visual information such as actions and enriches the descriptions using a speech recognition model. Though using expert models is straightforward, it may not be as flexible as adopting a learnable interface. The conversion of foreign modalities into text would cause information loss. For example, transforming videos into textual descriptions distorts spatial-temporal relationships [21].
@@ 3.1.1 Training detail
As the first training stage, pre-training mainly aims to align different modalities and learn multimodal world knowledge. Pre-training stage generally entails large-scale text-paired data, e.g. caption data. Typically, the caption pairs describe images/audio/videos in natural language.
Here, we consider a common scenario where MLLMs are trained to align vision with text. As illustrated in Table 3, given an image, the model is trained to predict autoregressively the caption of the image, following a standard cross-entropy loss. A common approach for pre-training is to freeze pre-trained modules (e.g. visual encoders and LLMs) and train a learnable interface [16]. The idea is to align different modalities without losing pre-trained knowledge. Some methods [28] also unfreeze more modules (e.g. visual encoder) to enable more trainable parameters for alignment. It should be noted that the training scheme is closely related to data quality. For short and noisy caption data, using lower resolution (e.g. 224) can speed up the training process, while for longer and cleaner data, it is better to utilize higher resolutions (e.g. 448 or higher) to mitigate hallucinations. Besides, ShareGPT4V [64] finds that with high-quality caption data in the pretraining stage, unlocking the vision encoder promotes better alignment.
@@
Table 3: A simplified template to structure the caption data. {<image>} is the placeholder for the visual tokens, and {caption} is the caption for the image. Note that only the part marked in red is used for loss calculation.

### v4

@@ Abstract
Recently, Multimodal Large Language Model (MLLM) represented by GPT-4V has been a new rising research hotspot, which uses powerful Large Language Models (LLMs) as a brain to perform multimodal tasks. The surprising emergent capabilities of MLLM, such as writing stories based on images and OCR-free math reasoning, are rare in traditional multimodal methods, suggesting a potential path to artificial general intelligence. To this end, both academia and industry have endeavored to develop MLLMs that can compete with or even better than GPT-4V, pushing the limit of research at a surprising speed. In this paper, we aim to trace and summarize the recent progress of MLLMs. First of all, we present the basic formulation of MLLM and delineate its related concepts, including architecture, training strategy and data, as well as evaluation. Then, we introduce research topics about how MLLMs can be extended to support more granularity, modalities, languages, and scenarios. We continue with multimodal hallucination and extended techniques, including Multimodal ICL (M-ICL), Multimodal CoT (M-CoT), and LLM-Aided Visual Reasoning (LAVR). To conclude the paper, we discuss existing challenges and point out promising research directions. In light of the fact that the era of MLLM has only just begun, we will keep updating this survey and hope it can inspire more research. An associated GitHub link collecting the latest papers is available at https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models.
Index Terms:
Multimodal Large Language Model, Vision Language Model, Large Language Model.
@@ 1 Introduction
Recent years have seen the remarkable progress of LLMs [1, 2, 3, 4, 5]. By scaling up data size and model size, these LLMs raise extraordinary emergent abilities, typically including instruction following  [5, 6], In-Context Learning (ICL)  [7], and Chain of Thought (CoT)  [8]. Although LLMs have demonstrated surprising zero/few-shot reasoning performance on most Natural Language Processing (NLP) tasks, they are inherently “blind” to vision since they can only understand discrete text. Concurrently, Large Vision Models (LVMs) can see clearly [9, 10, 11, 12], but commonly lag in reasoning.
@@
In light of this complementarity, LLM and LVM run towards each other, leading to the new field of Multimodal Large Language Model (MLLM). Formally, it refers to the LLM-based model with the ability to receive, reason, and output with multimodal information. Prior to MLLM, there have been a lot of works devoted to multimodality, which can be divided into discriminative [13, 14, 15] and generative [16, 17, 18] paradigms. CLIP [13], as a representative of the former, projects visual and textual information into a unified representation space, building a bridge for downstream multimodal tasks. In contrast, OFA [16] is a representative of the latter, which unifies multimodal tasks in a sequence-to-sequence manner. MLLM can be classified as the latter according to the sequence operation, but it manifests two representative traits compared with the traditional counterparts: (1) MLLM is based on LLM with billion-scale parameters, which is not available in previous models. (2) MLLM uses new training paradigms to unleash its full potential, such as using multimodal instruction tuning [19, 20] to encourage the model to follow new instructions. Armed with the two traits, MLLM exhibits new capabilities, such as writing website code based on images [21], understanding the deep meaning of a meme [22], and OCR-free math reasoning [23].
@@
Figure 1: A timeline of representative MLLMs. We are witnessing rapid growth in this field. More works can be found in our released GitHub page, which is updated daily.
@@
Ever since the release of GPT-4 [3], there has been a research frenzy over MLLMs because of the amazing multimodal examples it shows. Rapid development is fueled by efforts from both academia and industry. Preliminary research on MLLMs focuses on text content generation grounded in text prompts and image [20, 24]/video [25, 26]/audio [27]. Subsequent works have expanded the capabilities or the usage scenarios, including: (1) Better granularity support. Finer control on user prompts is developed to support specific regions through boxes [28] or a certain object through a click [29]. (2) Enhanced support on input and output modalities [30, 31], such as image, video, audio, and point cloud. Besides input, projects like NExT-GPT [32] further support output in different modalities. (3) Improved language support. Efforts have been made to extend the success of MLLMs to other languages (e.g. Chinese) with relatively limited training corpus [33, 34]. (4) Extension to more realms and usage scenarios. Some studies transfer the strong capabilities of MLLMs to other domains such as medical image understanding [35, 36, 37] and document parsing [38, 39, 40]. Moreover, multimodal agents are developed to assist in real-world interaction, e.g. embodied agents [41, 42] and GUI agents [43, 44, 45]. An MLLM timeline is illustrated in Fig. 1.
@@
The following parts of the survey are structured as such: the survey starts with a comprehensive review of the essential aspects of MLLMs, including (1) Mainstream architectures (§2); (2) A full recipe of training strategy and data (§3); (3) Common practices of performance evaluation (§4). Then, we delve into a deeper discussion on some important topics about MLLMs, each focusing on a main problem: (1) What aspects can be further improved or extended (§5)? (2) How to relieve the multimodal hallucination issue (§6)? The survey continues with the introduction of three key techniques (§7), each specialized in a specific scenario: M-ICL (§7.1) is an effective technique commonly used at the inference stage to boost few-shot performance. Another important technique is M-CoT (§7.2), which is typically used in complex reasoning tasks. Afterward, we delineate a general idea to develop LLM-based systems to solve composite reasoning tasks or to address common user queries (§7.3). Finally, we finish our survey with a summary and potential research directions.
@@ 2 Architecture
TABLE I: A summary of commonly used image encoders.
@@ 2.1 Modality encoder
The encoders compress raw information, such as images or audio, into a more compact representation. Rather than training from scratch, a common approach is to use a pre-trained encoder that has been aligned to other modalities. For example, CLIP [13] incorporates a visual encoder semantically aligned with the text through large-scale pre-training on image-text pairs. Therefore, it is easier to use such initially pre-aligned encoders to align with LLMs through alignment pre-training (see §3.1).
The series of commonly used image encoders are summarized in Table I. Apart from vanilla CLIP image encoders [13], some works also explore using other variants. For example, MiniGPT-4 [21] adopts an EVA-CLIP [48, 47] (ViT-G/14) encoder, which is trained with improved training techniques. In contrast, Osprey [29] introduces a convolution-based ConvNext-L encoder [46] to utilize higher resolution and multi-level features. Some works also explore encoder-free architecture. For instance, the image patches of Fuyu-8b [49] are directly projected before sending to LLMs. Thus, the model naturally supports flexible image resolution input.
When choosing encoders, one often considers factors like resolution, parameter size, and pretraining corpus. Notably, many works have empirically verified that using higher resolution can achieve remarkable performance gains [34, 50, 51, 52]. The approaches for scaling up input resolution can be categorized into direct scaling and patch-division methods. The direct scaling way inputs images of higher resolutions to the encoder, which often involves further tuning the encoder [34] or replacing a pre-trained encoder with higher resolution [50]. Similarly, CogAgent [44] uses a dual-encoder mechanism, where two encoders process high and low-resolution images, respectively. High-resolution features are injected into the low-resolution branch through cross-attention. Patch-division methods cut a high-resolution image into patches and reuse the low-resolution encoder. For example, Monkey [51] and SPHINX [53] divide a large image into smaller patches and send sub-images together with a downsampled high-resolution image to the image encoder, where the sub-images and the low-resolution image capture local and global features, respectively. In contrast, parameter size and training data composition are of less importance compared with input resolution, found by empirical studies [52].
Similar encoders are also available for other modalities. For example, Pengi [27] uses CLAP [54] model as the audio encoder. ImageBind-LLM [30] uses the ImageBind [55] encoder, which supports encoding image, text, audio, depth, thermal, and Inertial Measurement Unit (IMU) data. Equipped with the strong encoder, ImageBind-LLM can respond to the input of multiple modalities.
TABLE II: A summary of commonly used open-sourced LLMs. en, zh, fr, and de stand for English, Chinese, French, and German, respectively.
@@ 2.2 Pre-trained LLM
We summarize the commonly used and publicly available LLMs in Table II. Notably, most LLMs fall in the causal decoder category, following GPT-3 [7]. Among them, Flan-T5 [56] series are relatively early LLMs used in works like BLIP-2 [59] and InstructBLIP [60]. LLaMA series [5, 57] and Vicuna family [4] are representative open-sourced LLMs that have attracted much academic attention. Since the two LLMs are predominantly pre-trained on English corpus, they are limited in multi-language support, such as Chinese. In contrast, Qwen [58] is a bilingual LLM that supports Chinese and English well.
It should be noted that scaling up the parameter size of LLMs also brings additional gains, similar to the case of increasing input resolution. Specifically, Liu et al. [50, 61] find that simply scaling up LLM from 7B to 13B brings comprehensive improvement on various benchmarks. Furthermore, when using a 34B LLM, the model shows emergent zero-shot Chinese capability, given that only English multimodal data are used during training. Lu et al. [62] see a similar phenomenon by scaling up LLMs from 13B to 35B and 65B/70B, where the larger model size brings consistent gains on benchmarks specifically designed for MLLMs. There are also works that use smaller LLMs to facilitate deployment on mobile devices. For example, MobileVLM series [63, 64] use downscaled LLaMA [5] (termed as MobileLLaMA 1.4B/2.7B), enabling efficient inference on mobile processors.
Recently, explorations of Mixture of Experts (MoE) architecture for LLMs have garnered rising attention [65, 66, 67]. Compared with dense models, the sparse architecture enables scaling up total parameter size without increasing computational cost, by selective activation of the parameters. Empirically, MM1 [52] and MoE-LLaVA [68] find that MoE implementation achieves better performance than the dense counterpart on almost all the benchmarks.
@@ 2.3 Modality interface
Since LLMs can only perceive text, bridging the gap between natural language and other modalities is necessary. However, it would be costly to train a large multimodal model in an end-to-end manner. A more practical way is to introduce a learnable connector between the pre-trained visual encoder and LLM. The other approach is to translate images into languages with the help of expert models, and then send the language to LLM.
Learnable Connector. It is responsible for bridging the gap between different modalities. Specifically, the module projects information into the space that LLM can understand efficiently. Based on how multimodal information is fused, there are broadly two ways to implement such interfaces, i.e. token-level and feature-level fusion.
For token-level fusion, features output from encoders are transformed into tokens and concatenated with text tokens before being sent into LLMs. A common and feasible solution is to leverage a group of learnable query tokens to extract information in a query-based manner [69], which first has been implemented in BLIP-2 [59], and subsequently inherited by a variety of work [70, 60, 26]. Such Q-Former-style approaches compress visual tokens into a smaller number of representation vectors. In contrast, some methods simply use a MLP-based interface to bridge the modality gap [20, 37, 71, 72]. For example, LLaVA series adopts one/two linear MLP [20, 50] to project visual tokens and align the feature dimension with word embeddings.
@@
As another line, feature-level fusion inserts extra modules that enable deep interaction and fusion between text features and visual features. For example, Flamingo [74] inserts extra cross-attention layers between frozen Transformer layers of LLMs, thereby augmenting language features with external visual cues. Similarly, CogVLM [75] plugs in a visual expert module in each Transformer layer to enable dual interaction and fusion between vision and language features. For better performance, the QKV weight matrix of the introduced module is initialized from the pre-trained LLM. Similarly, LLaMA-Adapter [76] introduces learnable prompts into Transformer layers. These prompts are first embedded with visual knowledge and then concatenated with text features as prefixes.
@@
Expert Model. Apart from the learnable interface, using expert models, such as an image captioning model, is also a feasible way to bridge the modality gap [77, 78, 79, 80]. The basic idea is to convert multimodal inputs into languages without training. In this way, LLMs can understand multimodality by the converted languages. For example, VideoChat-Text [25] uses pre-trained vision models to extract visual information such as actions and enriches the descriptions using a speech recognition model. Though using expert models is straightforward, it may not be as flexible as adopting a learnable interface. The conversion of foreign modalities into text would cause information loss. For example, transforming videos into textual descriptions distorts spatial-temporal relationships [25].
@@ 3.1.1 Training Detail
As the first training stage, pre-training mainly aims to align different modalities and learn multimodal world knowledge. Pre-training stage generally entails large-scale text-paired data, e.g. caption data. Typically, the caption pairs describe images/audio/videos in natural language sentences.
Here, we consider a common scenario where MLLMs are trained to align vision with text. As illustrated in Table III, given an image, the model is trained to predict autoregressively the caption of the image, following a standard cross-entropy loss. A common approach for pre-training is to keep pre-trained modules (e.g. visual encoders and LLMs) frozen and train a learnable interface [72, 20, 35]. The idea is to align different modalities without losing pre-trained knowledge. Some methods [81, 82, 34] also unfreeze more modules (e.g. visual encoder) to enable more trainable parameters for alignment. It should be noted that the training scheme is closely related to the data quality. For short and noisy caption data, a lower resolution (e.g. 224) can be adopted to speed up the training process, while for longer and cleaner data, it is better to utilize higher resolutions (e.g. 448 or higher) to mitigate hallucinations. Besides, ShareGPT4V [83] finds that with high-quality caption data in the pretraining stage, unlocking the vision encode promotes better alignment.
@@
TABLE III: A simplified template to structure the caption data. {<image>} is the placeholder for the visual tokens, and {caption} is the caption for the image. Note that only the part marked in red is used for loss calculation.